Gomi Waste Disposal is evaluating a project that would require the purchase of a piece of equipment for 115,000 dollars today. During year 1, the project is expected to have relevant revenue of 116,000 dollars, relevant costs of 32,000 dollars, and relevant depreciation of 28,000 dollars. Gomi Waste Disposal would need to borrow 115,000 dollars today to pay for the equipment and would need to make an interest payment of 3,000 dollars to the bank in 1 year. Relevant net income for the project in year 1 is expected to be 32,150 dollars. What is the tax rate expected to be in year 1?
Answer #1

Year 1:

Let the expected tax rate be t

Relevant Revenue =$ 116000

Less: Relevant Costs = $ 32000

EBITDA = $ 84000

Less: Depreciation Expense = $ 28000

EBIT = $ 56000

Less: Interest Payment = $ 3000

Profit Before Tax = $ 53000

Less: Tax Expense = 53000 x t

Net Income = 53000 - 53000 x t = 53000 x (1-t) = $ 31250

(1-t) = (31250/53000) = 0.6066

t = 0.3934 or 39.34 %